OpenStudy (juan1857):

What are the points of discontinuity? Are they all removable? y=(x+4)/(x^2+5x+4)

OpenStudy (johnweldon1993):

Hint* think about how you can factor the denominator

OpenStudy (juan1857):

(x+4)(x+1)

OpenStudy (welshfella):

so now we can simplify the function

OpenStudy (juan1857):

so would it be 1/(x+1)?

OpenStudy (welshfella):

right so where would there be a discontinuity in the graph of this function?

OpenStudy (juan1857):

-1?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

it looked like there was another one at x = -4 but since the (x + 4) cancelled out its was removed.

OpenStudy (welshfella):

there's a discontinuity at x = -1

OpenStudy (juan1857):

ok, so how do you know if it is removable?

OpenStudy (welshfella):

https://www.desmos.com/calculator/o40hkz9gd

OpenStudy (welshfella):

I guess because we could remove it!

OpenStudy (juan1857):

so is that the only point of discontinuity?

OpenStudy (welshfella):

to be honest I haven't come across that expression 'removable discontinuity' before , that's why im sort of guessing....

OpenStudy (juan1857):

oh ok

OpenStudy (welshfella):

Yes just one point of discontimuity

OpenStudy (welshfella):

as you see Desmos gives that as the graph of the original function

OpenStudy (juan1857):

alright, Thanks

OpenStudy (welshfella):

yw

OpenStudy (juan1857):

still need help with finding if it is removable

OpenStudy (marcelie):

hmm i think it would be removal at x= -4 ?

satellite73 (satellite73):

yes, because when you factor and cancel you are removing it

OpenStudy (juan1857):

oh so 1, isn't a removable?

satellite73 (satellite73):

nope

satellite73 (satellite73):

the finished product is \[\frac{1}{x+1}\]so the discontinuity at \(-1\) is still there

OpenStudy (marcelie):

well ik this is not the right term to say it but whenever you see the same factor that cancels out that would be your removable

satellite73 (satellite73):

sounds right to me

OpenStudy (juan1857):

so you always get the opposite sign ?

OpenStudy (marcelie):

oh okay x) its been awhile havent seen this lol

satellite73 (satellite73):

\[\frac{x+4}{(x+1)(x+4)}\] now you see it (the discontinuity at -4 \[=\frac{1}{x+1}\] now you don't

satellite73 (satellite73):

as for the "opposite sign" when you set \[x+1=0\] and solve you get \(x=-1\)

OpenStudy (juan1857):

so for the removable you get the opposite sign too?

OpenStudy (juan1857):

because it was (x+4) so it should be -4, to make it 0, right?

OpenStudy (juan1857):

because it was (x+4) so it should be -4, to make it 0, right?

OpenStudy (marcelie):